### v1

@@ 1 Introduction
In the past decade, large direct imaging campaigns using extreme adaptive optics (extreme AO) systems coupled with near-infrared (IR) integral field spectrographs have discovered jovian planets and brown dwarfs on solar system scales and characterized their atmospheres (e.g. Macintosh et al., 2015; Chauvin et al., 2017; Konopacky et al., 2016; Rajan et al., 2017; Cheetham et al., 2019; Currie et al., 2023b). Most of these programs choose targets in an unbiased way, selecting stars for imaging observations based on properties like age and distance (Nielsen et al., 2019; Vigan et al., 2021; Currie et al., 2023a). Unfortunately, despite surveying hundreds of stars, such “blind” surveys have resulted in a low yield of new substellar companion discoveries: e.g. two new companions out of 300 targeted with the Gemini Planet Imager Exoplanet Survey.
@@ 2.3 Radial-Velocity Data From the SOPHIE Archive
To our high-contrast imaging data, we add precision radial-velocity (RV) data from the SOPHIE instrument (Bouchy & Sophie Team, 2006; Bouchy et al., 2013; Perruchot et al., 2008). The data were taken between 2006 October 14 and 2018 December 6 and extend the RV coverage from Currie et al. (2020a), which was drawn entirely from the Lick Observatory archive data using the Hamilton spectrograph (Fischer et al., 2014). The spectra were taken in high-resolution mode using the Thorium-Argon lamp for wavelength determination. The RV drift caused by a spectrograph instability was monitored by simultaneously observing Thorium-Argon lamp or Fabry-Pérot étalon. Exposure times ranged between 120 and 1200 .
@@ 4 Dynamical Analysis
For Simulation 2 adding SOPHIE’s RVs data, we run orvara with a total of 25 temperatures, with 100 walkers for each temperature and 500,000 steps per walkers. The MCMC converged with a typical autocorrelation timescale of 45 steps per parameter. We display in Table A4 the fitted parameters and the priors of HD33632 Ab. Figure 10 displays posterior distributions of orbital parameters and the BD’s mass for this simulation.
@@ 5 Discussion
Our dynamical analysis illustrates the key role that complementary RV data play even for companion search programs focused on direct imaging and astrometric data. In contrast with earlier HD 33632 results comparing modeling results with single-epoch imaging data vs. two epochs (Currie et al., 2021), we found little improvement in mass and orbit constraints on HD 33632 Ab from additional positional measurements from imaging777It is possible that this result is specific to only a subset of brown dwarfs and planets detected from imaging and astrometry. Analysis of other substellar companions detected with imaging and astrometry yields progressively tighter constraints on mass and orbital properties as additional imaging data are accumulated (T.C., unpublished). . However, when new RV data were incorporating into our modeling, the width of the posterior distributions for HD 33632 Ab’s mass and orbital properties generally shrunk by about a factor of 1.5 to 2 (e.g. a mass of ). RV data offer a fuller representation of companion’s dynamical influence on its host star when combined with astrometry. They should be obtained when possible for other accelerating stars that are targets of planet and brown dwarf searches, especially Sun-like stars amenable to precision RV measurements (e.g. HIP 21152; Kuzuhara et al., 2022; Franson et al., 2023).
@@
Our study provides a complementary probe of HD 33632 Ab’s atmosphere using different data sources than Hsu et al. (2024). They focused on high-dispersion coronagraphic spectroscopy to detect water and carbon monoxide at peak SNRs of 4.8 and 2.7 using the cloudless Sonora Bobcat atmosphere models as molecular templates (Marley et al., 2021) and SNRs of 7.8 and 6.1 using the BT-Settl models and the their baseline forward retrieval model using the petitRadTrans code (Mollière et al., 2019), which both incorporate clouds. Like Hsu et al. (2024), we find evidence for disequilibrium chemistry in HD 33632 Ab’s atmosphere. Temperatures derived from their high-resolution data alone (1550–2500 ) imply spectral types of M6–L4 and are easily ruled out by both our empirical comparisons and atmospheric modeling analysis. However, their retrieval posteriors for temperature incorporating photometry and spectra from Currie et al. (2020a) are in better agreement with our results. Orbital parameters from this work are generally consistent with previous studies (Currie et al., 2020a; Brandt et al., 2021a). Our dynamical mass measurements without RVs are in slight tension with theirs (37 ) and our masses derived with the new SOPHIE RV data rule out their results at the 1.5 level. On the other hand, the usage of Gaia eDR3 data resolves the discrepancy between our parallax posterior distribution and theirs.

### v2

@@ 1 Introduction
In the past decade, large direct imaging campaigns using extreme adaptive optics (extreme AO) systems coupled with near-infrared (IR) integral field spectrographs have discovered jovian planets and brown dwarfs on solar system scales and characterized their atmospheres (e.g. Macintosh et al., 2015; Chauvin et al., 2017; Konopacky et al., 2016; Rajan et al., 2017; Cheetham et al., 2019; Currie et al., 2023b). Most of these programs select stars for imaging observations based on properties like age and distance without any direct dynamical evidence for a companion (Nielsen et al., 2019; Vigan et al., 2021; Currie et al., 2023a). Unfortunately, despite surveying hundreds of stars, such “blind” surveys have resulted in a low yield of new substellar companion discoveries: e.g. two new companions out of 300 targeted with the Gemini Planet Imager Exoplanet Survey.
@@ 2.3 Radial-Velocity Data From the SOPHIE Archive
To our high-contrast imaging data, we add precision radial-velocity (RV) data from the SOPHIE instrument (Bouchy & Sophie Team, 2006; Perruchot et al., 2008) and its upgraded version SOPHIE+ (Bouchy et al., 2013). The data were taken between 2006 October 14 and 2018 December 6 and extend the RV coverage (1998-1-18 to 2009-2-1) from Currie et al. (2020a), which was drawn entirely from the Lick Observatory archive data using the Hamilton spectrograph (Fischer et al., 2014). The spectra were taken in high-resolution mode using the Thorium-Argon lamp for wavelength determination. The RV drift caused by a spectrograph instability was monitored by simultaneously observing Thorium-Argon lamp or Fabry-Pérot étalon. We assumed that an offset between the RVs taken by SOPHIE and SOPHIE+ is insignificant based on the previous studies performed with those instruments (e.g., Bouchy et al., 2013; Heidari et al., 2024; Díaz et al., 2016; Kiefer et al., 2019). Exposure times ranged between 120 and 1200 .
@@ 4 Dynamical Analysis
For Simulation 2 adding SOPHIE’s RVs data, we run orvara with a total of 25 temperatures, with 100 walkers for each temperature and 500,000 steps per walkers. The MCMC converged with a typical autocorrelation timescale of 45 steps per parameter. Temperature is here is a concept in MCMC simulations with parallel tempering. MCMC chains run at different temperatures. Hot chains (high temperatures) allow for extended exploration of the parameter space, while cold chains (lower temperatures) constrain exploration to likelihood regions. The parallel tempering approach helps avoid local minima through state swaps between hot and cold chains, enabling the convergence of the MCMC run in the case of multi-mode distributions.
We display in Table A4 the fitted parameters and the priors of HD33632 Ab. Figure 10 displays posterior distributions of orbital parameters and the BD’s mass for this simulation.
@@ 5.1 Implications of Our Results For The Atmosphere of HD 33632 Ab
Formally, our results show some tension between dynamical masses and those implied from atmospheric modeling. A log(g) = 5, 0.97 RJup object has a mass of 38 , which is 20% lower than and inconsistent with our dynamical mass constraints.
These inconsistencies result from the limited grid phase space available for the Exo-REM models. Holding the radius at 0.97 , a surface gravity of log(g) = 5.15 – 0.15 dex higher than our best fit – would yield perfect agreement with our dynamical mass. Similarly, for a 1–2 Gyr-old system (see Brandt et al. (2021b)), the evolutionary models from Baraffe et al. (2003) predict surface gravities of 5.10–5.26. However, the publicly-available Exo-REM grid spans a slightly different range in surface gravity – log(g) = – with a coarse sampling of 0.5 dex. In other words, we cannot explore the model phase space that would bring our dynamical masses and atmospheric model-inferred masses into agreement. An Exo-REM grid extended to higher surface gravities may resolve this minor discrepancy.
Disequilibrium chemistry is an important contributor to consider in atmospheric models. Previous studies had shown that its effects are most significant for young, low-gravity objects near the L/T transition. However, for an older object, like HD 33632 Ab, weaker signatures of disequilibrium chemistry are displayed, which coarsely aligns with the spectrum. For example, its K-band spectrum contrasts with the L/T transition object HR 8799 d, whose rising slope in Fν space displays the absence of methane absorption. This difference suggests that HD 33632 Ab’s atmosphere is in a chemical equilibrium. Nevertheless, comparisons with Lacy & Burrows (2020) show that disequilibrium chemistry is still present to some extent, as equilibrium models alone overestimate the methane absorption feature.
5.2 The Importance of RV Data for Constraining the Orbit and Dynamical Mass of HD 33632 Ab
@@
Our dynamical analysis illustrates the key role that complementary RV data play even for companion search programs focused on direct imaging and astrometric data. In contrast with earlier HD 33632 results comparing modeling results with single-epoch imaging data vs. two epochs (Currie et al., 2021), we found little improvement in mass and orbit constraints on HD 33632 Ab from additional positional measurements from imaging777It is possible that this result is specific to only a subset of brown dwarfs and planets detected from imaging and astrometry. Analysis of other substellar companions detected with imaging and astrometry yields progressively tighter constraints on mass and orbital properties as additional imaging data are accumulated (T.C., unpublished). .
However, when new RV data were incorporating into our modeling, the width of the posterior distributions for HD 33632 Ab’s mass and orbital properties generally shrunk by about a factor of 1.5 to 2 (e.g. a mass of ). RV data offer a fuller representation of companion’s dynamical influence on its host star when combined with astrometry. They should be obtained when possible for other accelerating stars that are targets of planet and brown dwarf searches, especially Sun-like stars amenable to precision RV measurements (e.g. HIP 21152; Kuzuhara et al., 2022; Franson et al., 2023).
5.3 Comparisons to Other Recent HD 33632 Ab Studies
@@
Our study provides a complementary probe of HD 33632 Ab’s atmosphere using different data sources than Hsu et al. (2024). They focused on high-dispersion coronagraphic spectroscopy to detect water and carbon monoxide at peak SNRs of 4.8 and 2.7 using the cloudless Sonora Bobcat atmosphere models as molecular templates (Marley et al., 2021) and SNRs of 7.8 and 6.1 using the BT-Settl models and the their baseline forward retrieval model using the petitRadTrans code (Mollière et al., 2019), which both incorporate clouds. Like Hsu et al. (2024), we find evidence for disequilibrium chemistry in HD 33632 Ab’s atmosphere.
Temperatures derived from their high-resolution data alone (1550–2500 ) imply spectral types of M6–L4 and are easily ruled out by both our empirical comparisons and atmospheric modeling analysis. However, their retrieval posteriors for temperature incorporating photometry and spectra from Currie et al. (2020a) are in better agreement with our results. Orbital parameters from this work are generally consistent with previous studies (Currie et al., 2020a; Brandt et al., 2021a). Our dynamical mass measurements without RVs are in slight tension with theirs (37 ) and our masses derived with the new SOPHIE RV data rule out their results at the 1.5 level. On the other hand, the usage of Gaia eDR3 data resolves the discrepancy between our parallax posterior distribution and theirs.
Our work provides a contrast to Gibbs et al. (2024) who also investigated the dynamics and atmosphere of HD 33632 Ab primarily with SCExAO/CHARIS and band spectroscopy. Using previously-published RV data from Currie et al. (2020a), astrometry from Currie et al. (2020a), and a new CHARIS astrometric data point from October 2022, they estimate a dynamical mass of 51.6 : nearly identical to prior results from Brandt et al. (2021b). Our analysis – adding relative RV data from Hsu et al. (2024), including new RV data for the star, and expanding HD 33632 Ab’s astrometric baseline – yields consistent results but improves the precision on HD 33632 Ab’s mass, semimajor axis, eccentricity, and inclination by a factor of 2.
By fitting the CHARIS and band spectra individually to the ATMO-2020 and Sonora-Bobcat cloudless atmosphere models (Phillips et al., 2020; Marley et al., 2021), Gibbs et al. (2024) favor a temperature of 1700 . However, given HD 33632 Ab’s luminosity of log10() -4.62, their implied radius is 0.56 : inconsistent with major luminosity evolution predictions of 1 (e.g. Baraffe et al., 2003), likely unphysical, and reminiscent of early inconsistencies between IR data for L/T transition exoplanets and atmosphere models that are insufficiently cloudy (e.g. Currie et al., 2011). They are also inconsistent with expected temperatures for L/T transition objects constrained in other studies (e.g. Stephens et al., 2009). These shortcomings are likely due to the decision to model the CHARIS and band spectra individually instead of jointly and the neglect of the NIRC2 photometric data in their modeling.
5.4 Future Observations of HD 33632 Ab